OpenStudy (anonymous):

Show that tan(a-b)+tanb/1-tan(a-b)tanb=tana Please someone help me with this...

OpenStudy (anonymous):

this needs rewriting. hard to tell what the fractions are

OpenStudy (anonymous):

numerator is: tan(a-b)+tanb denominator is: 1-tan(a-b)tanb the whole thing equals to tana

OpenStudy (anonymous):

\[\tan(a)=\frac{ \tan(a-b)+\tan(b) }{ 1-\tan(a-b)\tan(b) }\]

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Use Addition Formula: \(tan(a+b) = \frac{tana+tanb}{1-tana.tanb}\)

OpenStudy (anonymous):

that's what I tried then I got:

OpenStudy (anonymous):

\[\tan(a)=\tan(a+b)(1-\tan(a)\tan(b))-\tan(b)\]

OpenStudy (anonymous):

\(\frac{tan(a-b) + tanb}{1-tan(a-b).tanb}=tan[(a-b)-b]=tana\)

OpenStudy (anonymous):

how do you get it to equal \[\tan[(a-b)-b]=tana\]

OpenStudy (neer2890):

tan[(a-b)+b]=tan a because tan (a+b)= (tan a + tan b)/(1- tan a tan b)

OpenStudy (anonymous):

but how do I get from the angle sum identity to the one with the double braquets?

OpenStudy (anonymous):

OMG I got big mistake

OpenStudy (anonymous):

Sorry @Smilesong123

OpenStudy (anonymous):

\(tan(a-b)+b=tana\) :)

OpenStudy (anonymous):

You just use Addition Formula

OpenStudy (neer2890):

here we are using tan(a+b) formula and writing tan(a-b){ as in your question} as tan a else you can use this formula \[\tan(a-b)= \frac{ \tan(a)- \tan(b }{ 1+ \tan(a)\tan(b)}\] in your question and solve the equation, you will get the same answer

OpenStudy (anonymous):

but how do you get that answer?

OpenStudy (anonymous):

\[\frac{tan\ a+tan\ b}{1-tan\ a.tan\ b}=\frac{\frac{sin\ a}{cos\ a}+\frac{sin\ b}{cos\ b}}{1-\frac{sin\ a}{cos\ a}\frac{sin\ b}{cos\ b}}=\frac{\frac{sina\ cosb+cosa\ sinb}{cosa\ cosb}}{\frac{cosa\ cosb - sina\ sinb}{cosa\ cosb}}=\frac{sin(a+b)}{cos(a+b)}=tan(a+b)\]

OpenStudy (anonymous):

\(let\ a(in\ formulas) = a-b\\ \quad b(in\ formulas) = b\)

OpenStudy (anonymous):

thank you so much! now it makes more sense! :D

OpenStudy (anonymous):

^_^